A circle is tangent to the x axis and has its center at (8,-5). Give the equation of the circle in polynomial form.
Is the answer= x2 + y2 - 16x + 10y + 64 ?
Answer by lwsshak3(11628) 
 A circle is tangent to the x axis and has its center at (8,-5). Give the equation of the circle in polynomial form.
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius.
For given circle:
center: (8,-5)
radius=5
equation: (x-8)^2+(y+5)^2=25
equation in polynomial form:
x^2-16x+64+y^2+10y+25=25
x^2+y^2-16x+10y=0
|
|
|
| |